Ratio proficiency McDougal Printing, Inc., had sales totaling $43,000,000 in fiscal year 2015. Some ratios for the company are listed below. Use this information to determine the dollar values of various income statement and balance sheet accounts as requested. Assume a 365-day year McDougal Printing, Inc. Year Ended December 31, 2015 Calculate values for the following: a. Gross profits b. Cost of goods sold c. Operating profits d. Operating expenses e. Earnings available for common stockholders f. Total assets g. Total common stock equity h. Accounts receivable les Gross profit margin Operating profit margin Net profit margin Return on total assets Return on common equity Total asset turnover Average collection period $43,000,000 82% 39% 8% 13.2% 19% 2.2 53.7 days
Answer #1
a Gross profits Sales×Gross profit margin $ 35,260,000 =43000000*82%
b Cost of goods sold Sales-Gross profit $    7,740,000 =43000000-35260000
c Operating profits Sales×Operating profit margin $ 16,770,000 =43000000*39%
d Operating expenses Gross profits-Operating profits $ 18,490,000 =35260000-16770000